Solve -11(8 – x) = x – 10(5 – x). Please show all of your work.
Answer by robertb(5830)

-88 + 11x = x - 50 + 10x
-88 + 11x = 11x - 50
-88 = -50, a contradiction.
Hence the solution set is the empty set.